OpenStudy (anonymous):

3^(2x) - 5(3^x) = -6

OpenStudy (anonymous):

\[3^{2x} - 5 ( 3^{x}) = -6\]

myininaya (myininaya):

try to let u=3^x so u^2=3^(2x) so try replace 3^x with u and 3^(2x) with u^2